

********************************************************************
UARS WEEKLY STATUS REPORT - 18 Sep 1998
********************************************************************
***********
SOLSTICE
***********
After ending daylight only operations on Monday (98257), SOLSTICE continues
to operate normally, collecting solar data on all available orbital tracking
periods and stellar calibration observations, whenever possible.
***********
SUSIM
***********
SUSIM restarted operations on Monday, Sept. 14. While operating,
SUSIM successfully gathered its normal solar, offset, and occultation
data during this past week without apparent problems.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Gordon and Marianna Shepherd are presenting papers on WINDII results
at the ICMA Working Group Workshop on "Layered Phenomena in the
Mesopause Region", being held in Kuehlungsborn, Germany, and at the
25th Annual European Meeting on "Atmospheric Studies by Optical Methods",
to be held in Granada, Spain, September 21-25.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
59 days HAL (4.80) PROD (L1) (1873R-1448C-33F=392L)
51 days HAL (4.80) PROD (L3) (1873R-827C=1046L)
74 days SOL (4.62) PROD (L2-L3) (2349R-1514C=835L)
16 days SUS (7.00) ENG (63R-57C=6L)
2 days WIN (5.10) ENG (4R-2C=2L)
The Fortran compiler has been upgraded from version 6.3 to 6.5 on CDHF2 and
CDHF5. A copy of version 6.3 was preserved and is available to the user
community for backward compatability testing, etc.
CDHF pending software upgrades: IDL 5.1, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 September 1998 (Orbit 38161, GMT Day 248) through 11
September 1998 (Orbit 38264, GMT Day 254).
MILESTONE
UARS passed seven (7) years in orbit on 12 September having been
transported to orbit by STS-48 from Kennedy Space Center at 7:11 pm on 12
September 1991. Eight of the ten instruments remain operational and
continue to successfully collect science data.
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 250/2235 38204 Off
SUSIM 250/2235 38204 Off
HRDI 250/2235 38204 Low Power
MLS 251/0527 38208 On
HALOE 254/2026 38263 Off
HALOE 254/2206 38264 On
The beta angle ranged between 39.0 degrees to 52.4 degrees for this report
period. The beta angle is now decreasing toward a minimum angle of -2.0
degrees for the yaw on 29 September 1998 (DOY 272).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.3 degrees C. Additional battery performance data
for each battery on 04 September 1998 and 11 September 1998 are provided in
Appendix D.
The clock error ranged from 10.3 to -10.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
¥ ALERTS = 00
¥ OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
¥ FOT 05
¥ ACRIM 01
¥ HALOE 02
¥ HRDI 01
¥ MLS 05
¥ SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
¥ FOT 01
Two (2) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/253 show the observatory data loss to be
21 hours, 05 minutes, 02.004 seconds (an increase of 0.480 seconds since
last report period). This is a 0.0345 percent data loss which equals a
99.9655 percent data capture for the mission.
The increase of 0.480 seconds occurred on Day 251 and was due to a hit in
data plus line errors.
The database modifications to the TCLKDLTA equation processor definition
successfully resolved the parameter type mismatch error that prevented
clock delta calculations from executing. The database was incorporated
with software version 10.12 on a new operational pack - PE49. The MSOCC
Pack compliment is now:
PACK # GND S/W VER DATBASE VER COMMENT
PE49 10.12 FOT 068 Prime
PE15 10.12 FOT 068 Backup
**** After PMON software uplink ****
PE51 10.12 FOT 069 Prime
PE03 10.12 FOT 069 Backup
The PE49 pack was tested and elevated to operational status in Orbit 38216
on Day 251.
The On Board Computer (OBC) Power Monitor (PMON) software update has been
rescheduled for 24 September 1998 at the request of the Project Scientist.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) No change to report.
User Planning System (UPS) - No change to report.
Multi Satellite Operations Control Center (MSOCC) - No change to report.
Command Management System (CMS) - Testing continues.
Flight Dynamics Facility (FDF) - Year 2K planning aids have been generated.
OBC PMON FLIGHT SOFTWARE UPDATE
Flight Software - Complete.
Ground Support Software - Complete.
Verification Testing - Delta time calculation problem isolated. Data base
update generated. New operational pack generated (PE49), tested, and
elevated to operations. Uplink delayed to 24 September 1998 due to
personnel availability.
IP NETWORK TESTING
Space Network (SN) - During the weekly scheduled IP support event, UARS was
unable to transmit commands from the POCC. UARS could not trouble shoot
the problem in R/T because of the playback length requirement. A switch to
the legacy lines was made to command the playback. All data was recovered.
Post pass investigation revealed a clock buffer overflow within the Multi-
Sat equipment and this contributed to the commanding problem.
Testing with the NCC 98 system to support routine flight operations is
being planned.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - Working with the DSN test lead for I/P testing to verify UARS
emergency backup sites.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The B/D thruster refill effort continues. The Eaton Latch Valve test
report is being reviewed by the UARS FOT to understand the variation of
valve open and close times with voltage; the duration between open and
close commands; and overlap of open command driving signal and close
command driving signal. This will allow creation of the command structure
(PDB's, RTS's, and Stored Commands) to satisfy the B/D hydrazine refill
requirements. The electronics are in manufacturing. Cables for the test
interconnection are waiting review before manufacturing. The Flight Remote
Interface Unit (RIU) is in Dennis Asata's possession.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing support to the re-engineering effort to
process the UARS telemetry on the TOMS/EP control center computer system.
UARS COMMAND RE-HOST PHASE 1 DEMONSTRATION
Flight operations continues to support the update of the Lockheed Martin
SCS-21 control center system to support a phase 1 feasibility
demonstration. The test script is being executed. Several problems have
been identified and solutions are being implemented.
MEETINGS
The next battery telecon is scheduled for 22 September 1998 at 1:00 pm.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 29
September 1998.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 September 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
38167 248/1202 98-324 OBC Clk Adj fr 20.6 to 0.00 msec/day
38193 250/0535 98-325 OBC Clk Adj fr 0.00 to 20.6 msec/day
38204 250/2235 98-326 ACRIM Power OFF
38204 250/2235 98-327 SUSIM Power OFF
38204 250/2235 98-328 HRDI Low Power Mode
38208 251/0527 98-329 MLS Power ON
38230 252/1700 98-330 MLS Clear Exception Monitor
38237 253/0315 98-331 OBC Clk Adj fr 20.6 to 0.00 msec/day
38240 253/0814 98-332 MLS Clear Exception Monitor
38248 253/2030 98-333 MLS MLS STOP/RUN(0)
38248 253/2030 98-334 OBC NOOP'd RTS Activation(s)
38258 254/1229 98-335 OBC Clk Adj fr 0.00 to 20.6 msec/day
38263 254/2026 98-336 HALOE Power OFF
38264 254/2206 98-337 MLS Scan Htr ON
38264 254/2206 98-338 HALOE Power ON
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 September 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 38161-69
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 38206-203,36-41,50-64
QTSISAIF Y-LO SSPP Baseplate by SISA I/F 38221-34
QTSSPPMX Y-HI SSPP -X Thermostat 38253,55,57
HATBAT Y-HI Thermistor Bolo. Area Temp. 38254,59-61
HRIFESURVTMP Y-HI IFE Survival Heater Temp. 38255
HRIFESURVTMP Y-HI HRDI IFE Survival Htr Temp 38073
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 September 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
38167 248/12:08 (05 Sept 98) from 20.6 to 00.0 msec/day
38193 250/05:40 (07 Sept 98) from 00.0 to 20.6 msec/day
38237 253/03:21 (10 Sept 98) from 20.6 to 00.0 msec/day
38258 254/12:35 (11 Sept 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 September 1998
04 Sep 1998 - GMT Day 247 Beta = 39.0 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.1 0.000 0.0
2 +39.2/-5.6 +6.44/+5.15 +28.8/-16.0 25.1 1.023 22.3
3 +44.8/-11.2 +5.47/+3.87 +28.4/-16.0 25.1 1.026 21.7
11 Sept 1998 - GMT Day 254 Beta = 52.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +39.2/-11.2 +5.79/+4.51 +26.0/-16.0 25.4 1.024 21.8
3 +44.8/-11.2 +4.51/+3.23 +25.2/-16.0 25.3 1.027 21.3
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 September 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-059 FOT 38244 253/1410 20186 Unable to command via IP
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-058 FOT 247/21:00 CMS LOAD TRANSFER
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-057 FOT Sept 98 See list 14170 GENERIC LATE ACQ
98-059 FOT 38224 253/14:10 20186 NO CMD/IP SUPPORT